OpenStudy (anonymous):

Solve the following differential equation: x ̈+3x ̇+2x=∂(t),x(0)=-1,x ̇(0)=2

OpenStudy (anonymous):

is that supposed to be: x''+3x'+2x=y''(x)? x'+3x+x^2+C=y'(x) x+3/2 x^2+ 1/3 x^3 +Cx +c=y(x) y(0)=-1 or y(-1)=0 ? y'(0)=2 or y(2)=0 ?

OpenStudy (nikvist):

∂(t) - what does this mean? Partial symbol or something other?